THE CYTOSINE WATER COMPLEX

A. M. DALY, S. MATA, \underline{C. BERMUDEZ}, M. BERDAKIN, I. PENA, C. CABEZAS, J. L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Edificio Quifima, Laboratorios de Espectroscopía y Bioespectroscopía, Parque Científico, Universidad de Valladolid, 47011 Valladolid, Spain.

A multi FID system has been adapted into the operation sequence^{a} of the LA-MB-FTMW spectrometer.^{b} Thanks to the reached sensitivity, one monohydrate of cytosine (A = 3725.61 (26) MHz, B = 980.385 (76) MHz, C = 777.231 (46) MHz) has been detected in the supersonic expansion.

aJ. –U. Grabow, W. Stahl, H. Dreizler, Rev. Sci. Instrum. 1996, 67, 4072 – 4084.

bJ. L. Alonso, C. Pérez, M. E. Sanz, J. C. López, S. Blanco, Phys. Chem. Chem. Phys. 2009, 11, 617 – 627.